QUESTION 19

  1. Kenny Silver, CFA, is estimating the price of a call option. The call has an exercise price of $100 and a remaining time to expiration of 273 days. The spot price of the underlying stock is $93.25 and a put of the same underlying stock, exercise price and remaining time to expiration is currently priced at $6.50. Assuming a risk-free rate of 8% and a 365-day period, the call option’s arbitrage-free price is

a.

$5.34

b.

$7.66

c.

$0 because the exercise price is greater than the spot price of the underlying stock.
Answer #1

As per Put Call Parity Equation,

C +(X)e-rt = P + St

C = 6.50 + 93.25 - (100)e-(0.08)(273/365)

C = $5.34

Option A is correct.